# Strict Optimality of Frequency Estimation Under Local Differential Privacy

### 12 Mar 2026 v1

@@ Abstract
This paper establishes the strict optimality in precision for frequency estimation under local differential privacy (LDP). We prove that a frequency estimator with a symmetric and extremal configuration, and a constant support size equal to an optimized value, is sufficient to achieve maximum precision. Furthermore, we derive that the communication cost of such an optimal estimator can be as low as $\log_{2}(\frac{d(d-1)}{2}+1)$, where $d$ denotes the dictionary size, and propose an algorithm to generate this optimal estimator.
In addition, we introduce a modified Count-Mean Sketch and demonstrate that it is practically indistinguishable from theoretical optimality with a sufficiently large dictionary size (e.g., $d=100$ for a privacy factor of $\epsilon=1$). We compare existing methods with our proposed optimal estimator to provide selection guidelines for practical deployment. Finally, the performance of these estimators is evaluated experimentally, showing that the empirical results are consistent with our theoretical derivations.
@@ 1 Introduction
Frequencies are among the most fundamental statistics of a dataset. They serve as the building blocks for deriving other statistical properties, ranging from mean and variance to higher-order moments, and are often used to infer the underlying probability distribution.
While computing frequencies is straightforward when an analyst has unrestricted access to a dataset, typically involving a simple iteration using a hash map, the task becomes significantly more complex when raw individual values are inaccessible. This is often the case when robust data privacy protections are implemented to mask sensitive information.
@@
LDP was proposed to mitigate these risks by perturbing data locally on the client side before it is sent to the server. Consequently, even if intercepted, an adversary cannot effectively infer the original value from the perturbed output. The degree of this perturbation is quantified by a privacy budget, $\epsilon$, representing the maximum amount of knowledge an adversary can gain from the localized result [pan2023knowledge].
Since frequency is such a fundamental statistical property, numerous LDP algorithms have been developed to estimate the frequencies within a dataset. These algorithms are discussed in detail in Section 1.1. Among them, the Subset Selection mechanism proposed by [wang2016mutual, ye2018subset] achieved state-of-the-art precision regarding $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses. These metrics, formally defined in Section 2.2, measure the overall precision across all estimated frequencies. Notably, since the introduction of Subset Selection, no subsequent LDP frequency estimator has surpassed its leading precision in these metrics.
Although the leading precision has remained unsurpassed for years, it remains unknown whether these algorithms are strictly optimal. The Subset Selection paper [ye2018subset] established a lower bound for LDP frequency estimation and proved that their algorithm is order-optimal when $e^{\epsilon}\ll d$, where $d$ denotes the dictionary size. However, when comparing the achieved leading precision with this derived lower bound, a significant gap remains in the constant term; specifically, the derived lower bound can be as low as $\frac{1}{512}$ of the achieved precision. It remains unclear whether an algorithm exists that can bridge this gap, or if existing mechanisms have already reached strict optimality without a formal proof.
This paper directly addresses these questions by establishing a strict lower bound for LDP frequency estimation precision in terms of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses:
@@
Denote $d$, $\epsilon$, $n$, and $\hat{f}$ as dictionary size (number of all possible inputs), privacy factor, dataset size, and frequency estimator, respectively. When $d\geq e^{\epsilon}+1,$ we have
$\min_{\hat{f}}\mathcal{L}_{1}(\hat{f})\approx\frac{\sqrt{2(d-1)[4de^{\epsilon}-(e^{\epsilon}+1)^{2}}]}{\pi\sqrt{n}(e^{\epsilon}-1)},$
$\min_{\hat{f}}\mathcal{L}_{2}(\hat{f})=\frac{(d-1)[4de^{\epsilon}-(e^{\epsilon}+1)^{2}]}{nd(e^{\epsilon}-1)^{2}}.$
Otherwise,
$\min_{\hat{f}}\mathcal{L}_{1}(\hat{f})\approx\frac{\sqrt{2d(d-1)(d+2e^{\epsilon}-2)}}{\pi\sqrt{n}(e^{\epsilon}-1)},$
$\min_{\hat{f}}\mathcal{L}_{2}(\hat{f})=\frac{(d-1)(d+2e^{\epsilon}-2)}{n(e^{\epsilon}-1)^{2}}.$
Full details are presented in Theorem 3.4 and Corollary 3.3.
@@
To establish these tight lower bounds, this paper first demonstrates the existence of an optimized frequency estimator built upon a uniformly random permutation of another frequency estimator, as detailed in Section 2.4. We then prove that the $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses for such an estimator are minimized when using an optimized reconstruction matrix and a constant support size for all responses from the LDP perturbation. Under these conditions, we express $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ as functions of the support size, assuming other parameters are given. Subsequently, we determine the optimal support size that minimizes $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, thereby obtaining their strictly optimized values. The complete proof is presented in Section 3.
@@
Furthermore, we derive that the communication cost, defined as the number of bits required to transfer a response from a client, for an optimized frequency estimator is bounded above by $\log_{2}(\frac{d(d-1)}{2}+1)$ where $d$ denotes the dictionary size. We then develop an optimized frequency estimator, Weighted Subset Selection, which achieves this reduced communication cost despite having high precomputation requirements. Subsequently, we evaluate the optimality of existing algorithms and find that Subset Selection [ye2018subset] already achieves strict optimality in terms of precision. We also demonstrate that the Count-Mean Sketch (CMS), with minor modifications, can approach strict optimality provided the dictionary size is sufficiently large, such as when $d=100$ for $\epsilon=1$. Notably, this modified version, which we refer to as Optimized CMS, requires a significantly lower communication cost that is logarithmic with respect to the dictionary size, making it much more efficient than Subset Selection. Consequently, we provide practical guidance for their deployment; the rule of thumb is to employ Optimized CMS when the dictionary size is sufficiently large and to use either Weighted Subset Selection or the original Subset Selection otherwise.
@@
Finally, we conduct two experiments in Section 6 to compare the performance of the optimized algorithms discussed in Section 5 with the theoretical strict lower bounds of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. Both experiments demonstrate that all three optimized algorithms align perfectly with the strict lower bounds.
@@ 1.1 Previous Work
Numerous LDP algorithms have been developed to estimate frequencies. Randomized Response [warner1965randomized, kairouz2014extremal], the first LDP algorithm existing even before the formal proposal of LDP, was designed to collect sensitive data and estimate their frequencies. However, the variance of its estimated frequencies increases linearly with the dictionary size, making it unsuitable for large dictionaries. To resolve this issue, RAPPOR [rappor] was proposed to estimate frequencies with precision independent of dictionary size, though its communication cost remains linear to the dictionary size. Subsequently, Hadamard Encoding [bassily2015hadamard] was developed to encode a client’s original value through the Hadamard matrix, reducing the communication cost to the logarithm of the dictionary size. Apple Inc. [apple_privacy] leveraged Hadamard Encoding to build a private Count Mean Sketch method. Local Hashing and asymmetric RAPPOR were invented by [wang2017localhashing] to achieve leading precision for estimating low frequencies. Meanwhile, Subset Selection was developed by [wang2016mutual, ye2018subset] with leading precision regarding $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ metrics. Later, [acharya2019hr] proposed Hadamard Response, which achieves precision identical to Subset Selection in high-privacy regimes ($\epsilon\rightarrow 0$). Subsequently, [chen2020rhr] improved Hadamard Response by proposing a recursive version; it offers similar (though slightly lower) precision compared to the leading benchmarks while maintaining a communication cost that is only logarithmic to the dictionary size. Projective Geometry was then developed by [feldman2022private], asymptotically achieving the same precision as Subset Selection [ye2018subset] with a logarithmic communication cost, although it requires specific dictionary sizes. Recently, [pan2025improving] improved the original Count Mean Sketch by [apple_privacy] and proved that it approaches the leading precision in $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ when the dictionary size is sufficiently large.
Several existing studies analyze the optimality of LDP frequency estimation. [kairouz2014extremal] demonstrated that any optimized LDP algorithm can be converted into an extremal configuration. Subsequently, [wang2017localhashing] proposed the symmetric configuration adopted in this paper and showed that all existing algorithms can be encompassed by this framework. While [ye2018subset] established lower bounds for $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses when $e^{\epsilon}\ll d$, these bounds are not strict and have constant terms that are significantly smaller than the strict bounds derived in this work.
@@ 2.1 Local Differential Privacy (LDP)
Local Differential Privacy (LDP) was initially proposed in [LDP] to ensure that querying an individual object does not significantly disclose its actual value. Consider an object $X\in[d]$ being queried by a mechanism $M$:
@@
$M$ is $\epsilon$-LDP if and only if $\forall x,x^{\prime}\in[d]$, and $\forall o\in\mathcal{O}=[m]$,
@@
where $\mathcal{O}$ represents the output space (all possible values) of $M$, and the space size is $m$. In this paper, we may abbreviate $(M(X)|X=x)$ as $M(x)$ when the specific object is not the primary focus and its value is denoted by $x$.
@@ 2.2 Frequency Estimation
Denote $X^{(i)}$ to be the $i$-th object of a dataset of size $n$, where each object takes a value from a dictionary $d]=\{1,2,..,d\}$. Thus, $\forall i\in[n]:X^{(i)}\in[d]$. The true frequency of a value $x$ is defined as $f(x)=\frac{\sum_{i\in[n]}\mathbbm{1}(X^{(i)}=x)}{n}$, where $\mathbbm{1}$ denotes an indicator function.
@@
    The server collects all responses $O^{(i)}]_{i\in[n]}$ and computes their frequency vector $\mathbf{f_{O}}=\frac{1}{n}\sum_{i=1}^{n}M(X^{(i)})$. The $o$-th element of this vector represents the frequency of the response $o$ across all responses.
@@
    The estimated frequency vector is computed as $\mathbf{\hat{f}}=Q\mathbf{f_{O}}$, where $Q$ is a $d\times m$ matrix (to be formally defined later). The $x$-th element, denoted $\hat{f}(x)$, serves as an estimate of the true frequency $f(x)$.
Let $P$ be the perturbation matrix of $M$, where $P_{o,x}=\Pr(M(x)=o)$, and let $\mathbf{f}$ be the true frequency vector such that its $x$-th element represents $f(x)$. Since $\mathbb{E}[\mathbf{f_{O}}]=P\mathbf{f}$, reconstructing $\mathbf{f}$ from $\mathbf{f_{O}}$ requires a linear transformation $Q$ such that $QP=I$, where $I$ is the identity matrix.
@@
$F(P,Q)=Q\mathbf{f_{O}}=\frac{1}{n}\sum_{i\in[n]}QM(X^{(i)})$ denotes a frequency estimator that uses $P$ to perturb and $Q$ to reconstruct
Given that $\mathbb{E}[Q\mathbf{f_{O}}]=QP\mathbf{f}=\mathbf{f}$ (assuming $QP=I$), we obtain the following result:
@@
$F(P,Q)$ is an unbiased frequency estimator.
@@
$\text{Var}(\hat{f})=(Q\circ Q)P\mathbf{f}-\mathbf{f},$
where $(Q\circ Q)$ denotes the point-wise (Hadamard) square, i.e., $(Q\circ Q)_{i,j}=Q_{i,j}^{2}$. The $x$-th element of this vector represents $\text{Var}(\hat{f}(x))$.

### 22 Mar 2026 v2

# Strict Optimality of Frequency and Distribution Estimation Under Local Differential Privacy

@@ Abstract
This paper establishes the strict optimality in precision for frequency and distribution estimation under local differential privacy (LDP). We prove that a linear estimator with a symmetric and extremal configuration, and a constant support size equal to an optimized value, is sufficient to achieve the theoretical lower bound of the $\mathcal{L}_{2}$ loss for both frequency and distribution estimation. The theoretical $\mathcal{L}_{1}$ lower bound is also achieved asymptotically. Furthermore, we derive that the communication cost of such an optimal estimator can be as low as $\log_{2}(\frac{d(d-1)}{2}+1)$ bits, where $d$ denotes the dictionary size, and propose an algorithm to generate this optimal estimator.
In addition, we introduce a modified Count-Mean Sketch and demonstrate that it is practically indistinguishable from theoretical optimality with a sufficiently large dictionary size (e.g., $d=100$ for a privacy parameter of $\epsilon=1$). We compare existing methods with our proposed optimal estimator to provide selection guidelines for practical deployment. Finally, the performance of these estimators is evaluated experimentally, showing that the empirical results are consistent with our theoretical derivations.
@@ 1 Introduction
The frequency of a dataset and the distribution of a population represent the most fundamental statistics. They serve as the essential building blocks for deriving a wide array of statistical properties, ranging from basic mean and variance to more complex higher-order moments.
While computing frequencies or distributions is straightforward when an analyst has unrestricted access to a dataset, typically involving a simple iteration using a hash map, the task becomes significantly more complex when raw individual values are inaccessible. This is often the case when robust data privacy protections are implemented to mask sensitive information.
@@
LDP was proposed to mitigate these risks by perturbing data locally on the client side before it is sent to the server. Consequently, even if intercepted, an adversary cannot effectively infer the original value from the perturbed output. The degree of this perturbation is quantified by a privacy parameter, $\epsilon$, representing the maximum amount of knowledge an adversary can gain from the perturbed result [pan2023knowledge].
Since frequency and distribution are such fundamental statistical properties, numerous LDP algorithms have been developed to estimate them effectively. These algorithms are discussed in detail in Section 1.1. Among them, the Subset Selection mechanism proposed by [wang2016mutual, ye2018subset] achieved state-of-the-art precision regarding $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses. These metrics, formally defined in Section 2.2, measure the overall precision across all estimated frequencies or distributions. Notably, since the introduction of Subset Selection, no subsequent LDP estimator has surpassed its leading precision in these metrics.
While the leading precision has remained unsurpassed for years, it remains unknown whether these algorithms are strictly optimal. The Subset Selection paper [ye2018subset] established a lower bound for LDP distribution estimation, proving that their algorithm is order-optimal when $e^{\epsilon}\ll d$. However, a significant gap remains in the constant term; specifically, the derived lower bound can be as low as $\frac{1}{512}$ of the achieved precision. A follow-up study [ye2019optimal] suggested that Subset Selection achieves asymptotic strict optimality for distribution estimation, but this does not apply to finite datasets or frequency estimation (see Section 1.1). It remains unclear whether an algorithm exists to bridge this gap, or if current approaches are already strictly optimal.
This paper directly addresses these questions by establishing a strict lower bound for LDP frequency and distribution estimation precision in terms of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses:
@@
Let $d$, $\epsilon$, $n$, and $\hat{f}$ denote the dictionary size (the number of all possible inputs), privacy parameter, dataset size, and frequency estimator, respectively. The minimum $\mathcal{L}_{2}$ loss is given by
$\min_{\hat{f}}\mathcal{L}_{2}(\hat{f})=\begin{cases}\displaystyle\frac{(d-1)[4de^{\epsilon}-(e^{\epsilon}+1)^{2}]}{nd(e^{\epsilon}-1)^{2}}&\text{if }d\geq e^{\epsilon}+1,\\[15.0pt] \displaystyle\frac{(d-1)(d+2e^{\epsilon}-2)}{n(e^{\epsilon}-1)^{2}}&\text{otherwise.}\end{cases}$
Meanwhile, $\min_{\hat{f}}\mathcal{L}_{1}(\hat{f})\approx\sqrt{\frac{2}{\pi}\min_{\hat{f}}\mathcal{L}_{2}(\hat{f})}$.
Full details are presented in Theorem 3.8 and Corollary 3.3. Distribution estimation has a similar result: $\min_{\hat{\theta}}\mathcal{L}_{2}(\hat{\theta},n)=\min_{\hat{f}}\mathcal{L}_{2}(\hat{f},n)+\frac{1-1/d}{n}$, and $\min_{\hat{\theta}}\mathcal{L}_{1}(\hat{\theta},n)\approx\sqrt{\frac{2}{\pi}\min_{\hat{\theta}}\mathcal{L}_{2}(\hat{\theta},n)}$, where $\hat{\theta}$ denotes a distribution estimator.
@@
To establish these tight lower bounds, this paper first demonstrates that an estimator constructed from a uniformly random permutation (URP), when paired with an optimized reconstruction matrix, achieves the theoretical $\mathcal{L}_{2}$ loss lower bound derived from the Fisher information for any given perturbation matrix. The theoretical $\mathcal{L}_{1}$ lower bound is also achieved asymptotically. We then prove that these losses are further minimized when the perturbation matrix maintains a constant support size across all responses. Under these conditions, we express $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ as functions of the support size while holding other parameters constant. Finally, we determine the optimal support size to derive the strict lower bounds for both loss functions. The complete proof is provided in Section 3.
@@
Furthermore, we derive that the communication cost, defined as the number of bits required to transfer a response from a client, for an optimal estimator is upper bounded by $\log_{2}(\frac{d(d-1)}{2}+1)$ bits, where $d$ denotes the dictionary size. We then develop an optimal estimator, Weighted Subset Selection, which achieves this reduced communication cost despite having high precomputation requirements. Subsequently, we evaluate the optimality of existing algorithms and find that Subset Selection [ye2018subset] already achieves strict optimality in terms of precision, while its communication cost is linear in $d$. We also demonstrate that the Count-Mean Sketch (CMS), with minor modifications, can approach strict optimality provided the dictionary size is sufficiently large, such as when $d=100$ for $\epsilon=1$. Notably, this modified version, which we refer to as Optimized CMS, requires a significantly lower communication cost of $O(\log_{2}(d))$ bits, making it much more efficient than Subset Selection. Consequently, we provide practical guidance for their deployment; the rule of thumb is to employ Optimized CMS when the dictionary size is sufficiently large and to use either Weighted Subset Selection or the original Subset Selection otherwise.
@@
Finally, we conduct two experiments in Section 6 to compare the performance of the optimal algorithms discussed in Section 5 with the theoretical strict lower bounds of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. Both experiments demonstrate that all three optimal algorithms align perfectly with the strict lower bounds.
@@ 1.1 Previous Work
Numerous LDP algorithms have been developed for frequency and distribution estimation. Randomized Response [warner1965randomized, kairouz2014extremal], predating the formal LDP definition, was initially designed for the frequency estimation of sensitive data. RAPPOR [rappor] later achieved estimation precision for individual values independent of the dictionary size $d$, although its communication cost remains linear in $d$. Subsequently, methods such as Hadamard Encoding [bassily2015hadamard], private Count-Mean Sketch [apple_privacy], Local Hashing and asymmetric RAPPOR[wang2017localhashing], Subset Selection [ye2018subset], Hadamard Response families [acharya2019hr, chen2020rhr], Projective Geometry [feldman2022private], Block Design [park2024exactly], Optimized Count-Mean Sketch [pan2025improving], and Modular Subset Selection [arcolezi2026private] were developed to enhance precision, with several reducing the communication cost to $O(\log d)$. Among these, Subset Selection [ye2018subset] achieves state-of-the-art precision for $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ metrics, while some algorithms [feldman2022private, pan2025improving] asymptotically approach this performance.
Several existing studies analyze the optimality of LDP distribution estimation. Kairouz et al. (2014) [kairouz2014extremal] demonstrated that any optimal LDP algorithm can be converted into an extremal configuration. While Ye and Barg (2018) [ye2018subset] established lower bounds for $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses when $e^{\epsilon}\ll d$, these bounds are not tight and have constant terms that are significantly smaller than the strict bounds derived in this work. A follow-up study [ye2019optimal] proved that Subset Selection achieves asymptotic strict optimality for $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ in distribution estimation. However, their proof relies on numerous approximations requiring an extremely large dataset size $n$. Furthermore, they do not address how Subset Selection deviates from optimal bounds when $n$ is finite. Finally, their proof of optimality for distribution estimation, which involves the probability mass function (PMF) of a multinomial distribution, does not directly translate to frequency estimation; the latter involves estimating a Poisson multinomial distribution, the PMF of which lacks a concise closed-form expression. In contrast, our paper derives the strict $\mathcal{L}_{2}$ lower bound for any dataset size in both frequency and distribution estimation, and provides a sufficient condition to achieve this bound.
@@ 2.1 Local Differential Privacy (LDP)
Local Differential Privacy (LDP) was initially proposed in [LDP] to ensure that querying an individual object does not significantly disclose its actual value. Consider an object $X\in[d]=\{1,2,..,d\}$ being queried by a mechanism $M$:
@@
$M$ is $\epsilon$-LDP if and only if $\forall x,x^{\prime}\in[d]$, and $\forall o\in[m]$,
@@
where $m$ represents the output domain size of $M$. In this paper, we may abbreviate $(M(X)|X=x)$ as $M(x)$ when the specific object is not the primary focus and its value is denoted by $x$.
@@ 2.2 Frequency Estimation
Denote $X^{(i)}$ to be the $i$-th object of a dataset of size $n$, where each object takes a value from a dictionary $[d]$. Thus, $\forall i\in[n]:X^{(i)}\in[d]$. The true frequency of a value $x$ is defined as $f(x)=\frac{\sum_{i\in[n]}\mathbbm{1}(X^{(i)}=x)}{n}$, where $\mathbbm{1}$ denotes an indicator function. Let $\mathbf{f}$ be the true frequency vector such that its $x$-th element represents $f(x)$.
@@
    The server collects all responses $O^{(i)}]_{i\in[n]}$ and computes their frequency vector $\mathbf{f_{O}}=\frac{1}{n}\sum_{i=1}^{n}M(X^{(i)})$. The $o$-th element of this vector represents the frequency of the output $o$ across all responses.
@@
    The estimated frequency vector is computed as $\mathbf{\hat{f}}=\hat{f}(\mathbf{f_{O}})$, where $\hat{f}$ is an estimator and the $x$-th element of $\mathbf{\hat{f}}$, denoted $\hat{f}(x)$, serves as an estimate of the true frequency $f(x)$.
To evaluate the overall precision of an estimator $\hat{f}$, we define $L_{1}$ and $L_{2}$ norms as:
$L_{1}(\hat{f},\mathbf{f},n)$ $=\mathbb{E}\left[\sum_{x=1}^{d}|\hat{f}(x)-f(x)|\right],$
$L_{2}(\hat{f},\mathbf{f},n)$ $=\mathbb{E}\left[\sum_{x=1}^{d}(\hat{f}(x)-f(x))^{2}\right].$
Assuming the estimator is unbiased (i.e., $\forall x:\mathbb{E}[\hat{f}(x)]=f(x)$), we have $\mathbb{E}[(\hat{f}(x)-f(x))^{2}]=\text{Var}(\hat{f}(x))$. Consequently, the $L_{2}$ norm can be expressed as:
$L_{2}(\hat{f},\mathbf{f},n)=\sum_{x=1}^{d}\text{Var}(\hat{f}(x)).$ (1)
Additionally, the worst-case norms (i.e., loss functions) are defined to be
$\mathcal{L}_{1|2}(\hat{f},n)=\max_{\mathbf{f}}L_{1|2}(\hat{f},\mathbf{f},n),$
where $\mathcal{L}_{1|2}$ can be either $\mathcal{L}_{1}$ or $\mathcal{L}_{2}$.
Let $P$ be the perturbation matrix of $M$, where $P_{o,x}=\Pr(M(x)=o)$. Since $\mathbb{E}[\mathbf{f_{O}}]=P\mathbf{f}$, a linear transformation $Q$ can reconstruct $\mathbf{f}$ from $\mathbf{f_{O}}$ if $QP=I$, where $I$ is the identity matrix.
@@
$\text{LN}(P,Q)=Q\mathbf{f_{O}}=\frac{1}{n}\sum_{i\in[n]}QM(X^{(i)})$ denotes a linear estimator that uses $P$ to perturb and $Q$ to reconstruct.
Given that $\mathbb{E}[Q\mathbf{f_{O}}]=QP\mathbf{f}=\mathbf{f}$, we obtain the following result:
@@
$\text{LN}(P,Q)$ is an unbiased frequency estimator.
@@
$\text{Var}(\hat{f})=\frac{1}{n}[(Q\circ Q)P\mathbf{f}-\mathbf{f}],$
where $(Q\circ Q)$ denotes the point-wise (Hadamard) square, i.e., $(Q\circ Q)_{i,j}=Q_{i,j}^{2}$.
@@
The relationship between $L_{1}$ and $\text{Var}(\hat{f}(x))$ is established in the following theorem:
Theorem 2.3.
The $L_{1}$ of $\hat{f}=\text{LN}(P,Q)$ is approximated as:
$\lim_{n\rightarrow\infty}L_{1}(\hat{f},\mathbf{f},n)=\sum_{x\in[d]}\sqrt{\frac{2}{\pi}\text{Var}(\hat{f}(x))}.$ (2)
The proof is provided in Appendix A.3.
2.3 Distribution Estimation
Assume that each object $X^{(i)}$ are independently sampled from an identical category distribution $\mathcal{C}(\bm{\theta})$, where the $x$-th element of $\bm{\theta}$ denotes the probability of generating value $x$. If the objects are perturbed by an LDP mechanism following Section 2.2, the estimator $\text{LN}(P,Q)$ also estimates the distribution $\bm{\theta}$. Denote $\hat{\bm{\theta}}$ as the estimated distribution, which satisfies
Theorem 2.4.
If $\hat{\theta}=\text{LN}(P,Q)$,
$\mathbb{E}[\hat{\theta}$] $=\bm{\theta}$